If a flexible container has an internal volume of 409 liters and gets heated from 92K to 19K what will the final pressure be in atmospheres (atm)? The initial pressure is 33 atm, and the final volume is 472 liters. Round your answer to the nearest whole number.
1
Expert's answer
2020-05-15T12:11:02-0400

P1 = 33 atm

V1 = 409 L

T1 = 92 K

P2 = unknown

V2 = 472 L

T2 = 19 K


Solution:

Combined Gas Law:

P1V1/T1 = P2V2/T2

or

P1V1T2 = P2V2T1

Thus:

(33 atm) × (409 L) × (19 K) = P2 × (472 L) × (92 K)

(256443 atm L K) = P2 × (43424 L K)

P2 = (256443 atm L K) / (43424 L K) = 5.906 atm = 6 atm

P2 = 6 atm


Answer: The final pressure will be 6 atm.